OpenStudy (ny,ny):

Solve for x: sin^2x + cos2x - cosx = 0 on the interval 0< x< 2pi

OpenStudy (ny,ny):

@nincompoop

OpenStudy (amorfide):

use your identity \[\sin^{2}x+\cos^{2}x=1\]

OpenStudy (ny,ny):

sin^2 x = 1 - cos^2x

OpenStudy (amorfide):

\[\cos(A+A)=cosAcosA-sinAsinA\]

OpenStudy (amorfide):

get them all in terms of cos then solve for x

OpenStudy (ny,ny):

isnt the formula cosA cosB - SinA sinB

OpenStudy (amorfide):

yes but you want cos2a which is cos(A+A) so where you have put B you put A

OpenStudy (amorfide):

well, replace a with x but you understand it right?

OpenStudy (ny,ny):

ok, i think so.

OpenStudy (amorfide):

if you have cos(A+B) but you want cos2A you let B=A because both angles are the same so you get Cos(A+A) as clear as I can explain it so you do the addition formula for that, then you will use your identity again to change your sin into terms of cos then you will substitute back into the original to solve for x

OpenStudy (amorfide):

give it a go let me know what you get

OpenStudy (ny,ny):

(1-cos^2x) + (cosx cosx - sinx sin x) - cos x = 0

OpenStudy (amorfide):

cosxcosx - sinxsinx = cos^2x - sin^2x

OpenStudy (amorfide):

so you are going to use the identity again

OpenStudy (ny,ny):

how did you get -sin^2x on the right side?

OpenStudy (amorfide):

\[1-\cos^{2}x+ (cosxcosx-sinxsinx)-cosx=0\] \[1-\cos^{2}x+ (\cos^{2}x-\sin^{2}x)-cosx=0\]

OpenStudy (ny,ny):

ok.

OpenStudy (amorfide):

use your identity again for sin squared then you can get a quadratic in cos squared then should be easy from there i gotta go to class

OpenStudy (ny,ny):

ok thanks

OpenStudy (amorfide):

@Kainui help him if he needs please

OpenStudy (ny,ny):

alrights.